# The Distribution of Atomic Hydrogen in the Host Galaxies of FRBs

### v2

@@ 2.2.1 GMRT Observations and Data Reduction
Through proposals 45059 and 46112, we obtained a total of 114 hours of band observation (85 hours on-source) centred on the hosts of FRB 20181220A, FRB 20181223C, FRB 20190418A, FRB 20190425A (chimefrbs) and FRB 20200223B (Ibik2023). These observations were taken in November 2023 and August 2024, using the GMRT Wideband Backend as the correlator. Depending on the location of the Hi line in the band, we used 100 MHz and 200 MHz bandwidths between 1260 - 1460 MHz, covering the entire range. These were subdivided into 4096 channels, resulting in velocity resolutions of km s-1 and 12 km s-1 respectively.
The raw data were transferred and processed using CASA (CASA) on the ilifu supercomputing cloud system. Initial flagging of corrupted antennas and scans was completed manually, and gain, bandpass, and phase calibrations were conducted following standard procedures. The data were rebinned to 1 MHz for self-calibration, and the calibration solutions were applied to the visibilities around the Hi line region. Final deep continuum images were made out to the first null of the uGMRT primary beam, whose resulting model continuum visibilities were subtracted from the corrected visibilities using the CASA task uvsub. The residual data were then run through a final round of automatic RFI removal.
A uv-plane continuum subtraction was completed with the CASA task uvcontsub, excluding 500 km s-1 around the region of the Hi line with a first order polynomial fit. We then made data cubes surrounding the host galaxy coordinates using the CASA task tclean with Briggs weighting. Various robust levels and resolutions were employed to test different emphases on sensitivity and structure, ensuring that around 5 pixels covered the beam. Due to some residual variations in the baseline, in a few cases we ran a final image-plane based continuum subtraction using the CASA task imcontsub, excluding the channels visibly containing Hi emission with a second order polynomial. Finally, each spectral-line datacube was convolved to its largest synthesised common beam to account for the frequency variation in the beam shape.
@@ 2.2.2 MeerKAT Observations and Data Reduction
Through three MeerKAT proposals (project IDs SCI-20220822-MG-01, SCI-20230907-MG-02 and SCI-20241103-MG-01), we obtained a total of 14 on-source hours, centred on the hosts of FRB 20210405I (frb20210405), FRB 20211127I (Glowacki_wallaby), FRB 20211212A, and FRB 20240201A (Shannon2024). The observations were conducted in December 2022, March 2023, January 2024, and February 2025 (capture block IDs 1670970986, 1679181682, and 1718998942). Each observation was taken in the band (107 MHz centred at 1358.5 MHz) at 32K and 32K-zoom spectral resolutions, which corresponds to velocity resolutions of 5.5 km s-1 and 0.7 km s-1 respectively.
@@
Model continuum visibilities were subtracted from the corrected data using uvsub, and then a further uvcontsub step was run. Spectral-line cubes were made using tclean with Briggs weighting, again with various robust levels, resolutions, and cleaning depths. All channels were convolved to a synthesised common beam, and then each cube was primary beam corrected using katbeam222https://github.com/ska-sa/katbeam.
@@ 2.2.3 FAST Observations and Data Reduction
The data are stored in FITS format as a dual polarisation time series of bandwidth data. We separated the on and off source data and performed a temperature calibration using the high power noise diode (e.g. Section 3.1 of Jing2023) to convert from raw power to brightness temperature. Due to the electromagnetic wave reflection that is common in single dish instruments, the FAST data contains periodic fluctuations in the bandpass which manifest as standing waves. We removed this signal using a fast Fourier transform (FFT) based approach, modelled off the methods of Xu2024. Using the spectroscopic redshifts of each target, we extracted 2000 km s-1 around the Hi line and stacked each observation weighted by its RMS noise.
@@ 3 Hi Results
FRB 20181220A | FAST | 0.72 | 9.419 | 315 2 | N/A | N/A | N/A | N/A |
GMRT | 0.76 | 9.443 | 126 2 | 2.72 | 2.04 | 20.5 10.0 | 11.0 | |
FRB 20181223C | FAST | 0.44 | 9.287 | 79 1 | N/A | N/A | N/A | N/A |
GMRT | 0.60 | 9.420 | 70 1 | 0.08 | 0.13 | 11.1 2.4 | 10.6 | |
FRB 20190418A | GMRT | 0.82 | 10.333 | 217 5 | 0.06 | 0.44 | 36.2 8.2 | - |
FRB 20190425A | FAST | 3.61 | 10.231 | 307 1 | N/A | N/A | N/A | N/A |
GMRT | 3.83 | 10.257 | 307 1 | 1.11 | 1.21 | 35.0 5.0 | 12.0 | |
FRB 20200223B | GMRT | 0.13 | 9.377 | N/A | 0.10 | 0.51 | N/A | N/A |
@@
FRB 20200723B | MeerKAT | 20.47 | 9.840 | 400 1 | 47.7 | 3.44 | 23.0 2.9 | 12.0 |
FRB 20201123A | MeerKAT | 0.41 | 9.722 | 415 10 | 5.15 | 9.53 | N/A | N/A |
FRB 20210405I | MeerKAT | 1.00 | 10.341 | 245 2 | 0.59 | 2.43 | 29.3 3.3 | 12.7 |
FRB 20211127I | MeerKAT | 0.90 | 9.99 | 158 1 | 1.16 | 2.46 | 14.5 2.5 | 11.7 |
FRB 20211212A | MeerKAT | 0.17 | 9.651 | 87 2 | 0.31 | 1.60 | 25.9 9.0 | - |
FRB 20231229A | Arecibo | 5.29 | 9.958 | 255 1 | N/A | N/A | N/A | N/A |
FRB 20231230A | MeerKAT | 0.18 | 8.891 | 206 6 | 5.19 | 3.98 | 12.3 3.4 | - |
FRB 20240201A | MeerKAT | 1.01 | 9.959 | 382 1 | 0.42 | 0.91 | - | - |
FAST | 0.89 | 9.901 | 384 1 | N/A | N/A | N/A | N/A | |
FRB 20250316A | Effelsberg | 7.13 | 9.127 | 132 2 | N/A | N/A | N/A | N/A |
@@
Estimates are accompanied by 10% (0.05 dex) uncertainties due to absolute flux scaling errors.
@@
Here we highlight the global Hi properties of the sample. For each spatially resolved observation (i.e., the GMRT and MeerKAT datacubes), we use the SoFiA-2 (Westmeier2021) source finding pipeline to extract Hi sources and generate moment maps. We use the pipeline’s default smooth-and-clip (S+C) finder with parameters following those used in Westmeier2022 for the pilot data release of the WALLABY survey (which SoFiA-2 was designed for). This includes a scfind.threshold value of 3.5 and a reliability.minSNR value of 3, to ensure only robust signals are picked up. For detections with particularly narrow line widths, we rerun SoFiA-2 with a maximum spectral axis smoothing kernel equal to this apparent width.
@@
Figure 1 presents spectra for each target, binned to lower spectral resolution where sensible. The GMRT and MeerKAT spectra are generated by summing the flux from pixels included in a 2D mask, which itself is a flattened projection of the 3D mask generated by SoFiA-2. We also present moment 0 (intensity), moment 1 (velocity), and moment 2 (velocity dispersion) maps for these targets in Figure 2, overlaid on imaging from the DESI Legacy Imaging Survey (DESI), the Dark Energy Camera Plane Survey (DECaPS, DECaPS), and the Pan-STARRs1 survey (PanSTARRS). The lowest contours are defined as the 3 level measured by SoFiA-2333See Table 5 in Appendix for column and surface densities..
@@
Table 3 presents a summary of the global Hi properties of the sample. These properties are measured from the masked spectra; i.e. only voxels in the 3D SoFiA-2 mask are included, and as a result they are slightly different to those displayed in Fig. 1. We calculate Hi masses following Meyer2017, assuming a 10% absolute flux scale uncertainty on all measurements due to calibration, which dominates the noise in all cases. As radio continuum (RC) emission is considered a robust tracer of star formation due to its association with young, massive stars, when RC sources are present, we estimate an SFR following the methods of Molnar2021.
@@
We estimate the radius of the Hi emission, HI, out to the 3 contour level by measuring the average surface density33footnotemark: 3 of the intensity maps as a function of elliptical radius from a chosen centre, which we choose to be the optical centre of brightness in the survey imaging. This HI estimate is not necessarily related to the disk size commonly used to describe HI disks e.g. for the mass-size relation (Wang2016), as many of our maps do not reach the nominal 1 pc-2 threshold used in the literature. We estimate the ellipticity and position angles of the galaxies using 2MASS (2MASS) K-band data where possible; otherwise we fit elliptical isophotes to the intensity map. These values are then combined with estimates of the galaxies’ rotation velocities to make broad approximations of the dynamical mass and hence the virial mass for these galaxies. We estimate using
@@
where is the line width at 50% of the peak flux and is the inclination of the galaxy; we discard three galaxies with inclinations lower than 10. To ensure these rotational velocities are in the flat part of each galaxies’ rotation curve, we examine position-velocity diagrams across the major axes and find that all targets extend beyond the break. We do not apply systematic corrections for instrumental broadening and turbulence as we expect such corrections to be small for the high resolution data, and we are not attempting a fully robust measurement. We calculate and following Yu2020.
@@ 3.1 Marginal / Non-Detections
As seen in Fig. 1, the host of FRB 20201123A peaks very marginally above the noise. This host is at a relatively high redshift of 0.0507, and the MeerKAT observation was 0.4 from the phase centre. Due to this, and the beam size, we presume the picked-up signal is simply the unresolved core emission; as such, we do not place much weight on its derived Hi properties, considering its Hi mass to be a lower limit. Only its intensity map is displayed, and we do not attempt to measure its asymmetry.
Fig. 1 also shows the non-detection of the host of FRB 20200223B. We place a 5 upper limit on its Hi mass using both the GMRT and FAST data, assuming a linewidth of 200 km s-1 and a disk size of 30 kpc. As outlined in Ibik2023, this host is in a transitional regime between star-formation and quiescence, which reflects its lower gas fraction. This is uncommon for FRB hosts; Gordon2023; Sharma2024 clearly find a preference towards star-forming galaxies, a conclusion well supported by our findings given that 13 of our 14 targets are emitting in Hi.
@@ 4.1 Asymmetry Metrics
There are a number of common metrics used in the literature to assess asymmetry (see the introduction of Deg2023, for an overview); in this study we utilise five: flux, spec, 1D, , and 3D. The first is a common historical measurement used to consider the lopsidedness of a spectrum, and is defined by
@@
The second metric, spec, was introduced by Reynolds2020LVHIS to emphasise more local disturbances through investigation of channel-by-channel variation, defined by
where and are the fluxes of the channel and its mirrored channel pair about a central channel. This centre is generally different than , and is instead chosen as the flux-weighted mean velocity to highlight the differences in profile far away from the centre of mass, e.g. through an extended tail on one side of the spectrum. These metrics were used in Michalowski21; Glowacki_wallaby to compare FRB hosts with wider LVHIS and HALOGAS Hi survey populations (Reynolds2020LVHIS), and thus we include calculations here.
Recently, Deg2023 introduced new metrics to properly account for the bias contributed by noise to the asymmetry of a profile, and to take advantage of the full information that three-dimensional datacubes can offer. As next generation surveys, such as WALLABY and LADUMA, will resolve many Hi galaxies, these higher dimensional metrics are likely to be the most robust way to quantify a galaxy’s disturbance, especially in low SNR regimes. Thus, we also calculate values for 1D, , and 3D, based on equation 20 in Deg2023:
@@
When extracting final values for each of the 1D metrics, we consider the optimal spectral resolution. Each can become temperamental as the SNR of the Hi line decreases and as the linewidth (in channels) decreases, and thus we take measurements at many different rebinning ratios to estimate the most stable resolution. This estimation differs between target and is done purely by eye; we inspect the variation of each metric to find the resolution which appears coincident with the most overall stable point in the three metrics, aiming to use the highest resolution possible. We highlight an example case in Fig. 13 in the Appendix.
@@ 4.2 Asymmetry Thresholds
We wish to identify the thresholds above which a profile is considered asymmetric for each of our metrics. Espada2011 modelled the flux distribution of AMIGA isolated galaxies, finding a 2 threshold of 1.26. Reynolds2020LVHIS did the same with the low density fields of LVHIS and HALOGAS, and the higher density field of VIVA, finding 2 thresholds of 1.25, 1.17, and 1.53 respectively for galaxies with stellar masses . Watts2020 focused on the more representative xGASS sample, and chose to define asymmetry by generating noiseless mock spectra with inherent flux equal to 1.1, perturbing them by various noise levels, and finding the threshold (as a function of SNR) within which 80% of the spectra are measured. Considering these results and the fact that a number of our targets are low SNR (10), we choose a flat threshold of 1.3; this is similar to the 3 limits of Espada2011; Reynolds2020LVHIS, and corresponds to the low SNR threshold of Watts2020.
Reynolds2020LVHIS also measured spec for the LVHIS, HALOGAS, and VIVA samples, finding 2 thresholds of 0.428, 0.263, and 2.031 respectively. As this metric is the most affected by noise, we choose a slightly arbitrary conservative threshold of 0.5; this also reflects the fact that FRB hosts are not generally found in low-density environments.
As the 1D, , and 3D measurements are even newer, there are not many comparisons to be made with wider populations from which to infer at what point a profile is considered asymmetric. Perron-Cormier2025 recently applied these metrics to data from the WALLABY pilot survey (Westmeier2022) and found that an 3D threshold of 0.5 corresponds to highly disturbed galaxies, though this value likely represents the extreme end of the disturbance spectrum. Fig. 8 of Deg2023 presents a trial analysis of 500 galaxies from a simulated WALLABY-like survey, showing a median visual disturbance of approximately 0.2, with most values falling within 0.1 of this median. Thus, an 3D value of 0.3 can serve as a rough threshold for identifying galaxies with above-average disturbance. In Section 5, we also present asymmetry measurements of the disturbed FRB host from Lee-waddell23 to use as a reference.

### v3

@@ 2.2.1 GMRT Observations and Data Reduction
Through proposals 45059 and 46112, we obtained a total of 114 hours of band observation (85 hours on-source) centred on the hosts of FRB 20181220A, FRB 20181223C, FRB 20190418A, FRB 20190425A chimefrbs (9) and FRB 20200223B Ibik2023 (43). These observations were taken in November 2023 and August 2024, using the GMRT Wideband Backend as the correlator. Depending on the location of the Hi line in the band, we used 100 MHz and 200 MHz bandwidths between 1260 - 1460 MHz, covering the entire range. These were subdivided into 4096 channels, resulting in velocity resolutions of km s-1 and 12 km s-1 respectively.
The raw data were transferred and processed using CASA CASA (62) on the ilifu supercomputing cloud system. Initial flagging of corrupted antennas and scans was completed manually, and gain, bandpass, and phase calibrations were conducted following standard procedures. The data were rebinned to 1 MHz for self-calibration, and the calibration solutions were applied to the visibilities around the Hi line region. Final deep continuum images were made out to the first null of the uGMRT primary beam, whose resulting model continuum visibilities were subtracted from the corrected visibilities using the CASA task uvsub. The residual data were then run through a final round of automatic RFI removal.
A uv-plane continuum subtraction was completed with the CASA task uvcontsub, excluding 500 km s-1 around the region of the Hi line with a first order polynomial fit. We then made data cubes surrounding the host galaxy coordinates using the CASA task tclean with Briggs weighting. Various robust levels and resolutions were employed to test different emphases on sensitivity and structure, ensuring that around 5 pixels covered the beam. Due to some residual variations in the baseline, in a few cases we ran a final image-plane based continuum subtraction using the CASA task imcontsub, excluding the channels visibly containing Hi emission with a second order polynomial. Finally, each spectral-line datacube was corrected for its frequency-dependent primary beam shape and convolved to its largest synthesised common beam.
@@ 2.2.2 MeerKAT Observations and Data Reduction
Through three MeerKAT proposals (project IDs SCI-20220822-MG-01, SCI-20230907-MG-02 and SCI-20241103-MG-01), we obtained a total of 14 on-source hours, centred on the hosts of seven FRBs. The observations were conducted from December 2022, through to August 2025 (capture block IDs 1670970986, 1679181682, 1718998942, 1739386395, 1740164485, 1751156931, 1753661382, and 1754260765). Each observation was taken in the band (107 MHz centred at 1358.5 MHz) at 32K and 32K-zoom spectral resolutions, which corresponds to velocity resolutions of 5.5 km s-1 and 0.7 km s-1 respectively.
@@
Model continuum visibilities were subtracted from the corrected data using uvsub, and then a further uvcontsub step was run. Spectral-line cubes were made using tclean with Briggs weighting, again with various robust levels, resolutions, and cleaning depths. Each cube was primary beam corrected using katbeam222https://github.com/ska-sa/katbeam and then was convolved to a synthesised common beam.
@@ 2.2.3 FAST Observations and Data Reduction
The data are stored in FITS format as a dual polarisation time series of bandwidth data. We separated the on and off source data and performed a temperature calibration using the high power noise diode (Jing2023, 45, e.g. Section 3.1 of) to convert from raw power to brightness temperature. Due to the electromagnetic wave reflection that is common in single dish instruments, the FAST data contains periodic fluctuations in the bandpass which manifest as standing waves. We removed this signal using a fast Fourier transform (FFT) based approach, modelled off the methods of Xu2024 (91). Using the spectroscopic redshifts of each target, we extracted 2000 km s-1 around the Hi line and stacked each observation weighted by its RMS noise.
@@ 3 Hi Results
| FRB 20181220A | GMRT | 0.76 0.12 | 9.44 0.08 | 126 14 | 2.72 | 2.29 | 20.5 10.0 | 11.0 |
| FAST | 0.72 0.08 | 9.42 0.05 | 315 2 | N/A | N/A | N/A | N/A | |
| FRB 20181223C | GMRT | 0.60 0.10 | 9.42 0.08 | 70 5 | 0.08 | 0.15 | 11.1 2.4 | 10.6 |
| FAST | 0.45 0.04 | 9.30 0.05 | 79 1 | N/A | N/A | N/A | N/A | |
| FRB 20190418A | GMRT | 0.82 0.15 | 10.33 0.09 | 217 19 | 0.06 | 0.51 | 36.2 8.2 | - |
| FRB 20190425A | GMRT | 3.55 0.39 | 10.22 0.05 | 303 3 | 1.11 | 1.35 | 35.0 5.0 | 12.0 |
| FAST | 3.61 0.36 | 10.23 0.05 | 307 1 | N/A | N/A | N/A | N/A | |
| FRB 20200223B | GMRT | 0.13 | 9.377 | N/A | 0.10 | 0.59 | N/A | N/A |
@@
| FRB 20200723B | MeerKAT | 20.47 2.11 | 9.84 0.05 | 401 4 | 47.7 | 3.42 | 23.0 2.9 | 12.0 |
| FRB 20201123A | MeerKAT | 0.41 0.18 | 9.72 0.16 | 415 36 | 5.15 | 10.9 | N/A | N/A |
| FRB 20210405I | MeerKAT | 1.00 0.12 | 10.34 0.06 | 245 5 | 0.59 | 2.85 | 29.3 3.3 | 12.7 |
| FRB 20211127I | MeerKAT | 0.90 0.10 | 9.99 0.05 | 158 3 | 1.16 | 2.82 | 14.5 2.5 | 11.7 |
| FRB 20211212A | MeerKAT | 0.17 0.04 | 9.65 0.11 | 87 8 | 0.31 | 1.92 | 25.9 9.0 | - |
| FRB 20231229A | Arecibo | 5.29 0.54 | 9.96 0.05 | 256 3 | N/A | N/A | N/A | N/A |
| FRB 20231230A | MeerKAT | 0.18 0.04 | 8.89 0.09 | 206 21 | 5.19 | 4.47 | 12.3 3.4 | - |
| FRB 20240201A | MeerKAT | 1.01 0.11 | 9.96 0.05 | 373 5 | 0.42 | 1.04 | - | - |
| FAST | 0.89 0.09 | 9.90 0.05 | 384 1 | N/A | N/A | N/A | N/A | |
| FRB 20240210A | MeerKAT | 1.51 0.16 | 9.61 0.05 | 166 1 | 1.67 | 1.19 | 16.8 4.2 | 11.38 |
| FRB 20240312D | MeerKAT | 1.25 0.13 | 10.19 0.05 | 182 1 | 1.26 | 3.31 | 27.2 4.3 | - |
| FRB 20240615B | MeerKAT | 0.25 0.03 | 9.84 0.06 | 76 1 | 0.30 | 1.99 | 30.3 3.0 | 11.39 |
| FRB 20250316A | Effelsberg | 7.13 0.79 | 9.13 0.05 | 132 2 | N/A | N/A | N/A | N/A |
@@
Estimates are accompanied by 10% uncertainties due to absolute flux scaling errors.
@@
Here we highlight the global Hi properties of the sample. For each spatially resolved observation (i.e., the GMRT and MeerKAT datacubes), we use the SoFiA-2 Westmeier2021 (89) source finding pipeline to extract Hi sources and generate moment maps. We use the pipeline’s default smooth-and-clip (S+C) finder with parameters following those used in Westmeier2022 (90) for the pilot data release of the WALLABY survey (which SoFiA-2 was designed for). This includes a scfind.threshold value of 3.5 and a reliability.minSNR value of 3, to ensure only robust signals are picked up. For detections with particularly narrow line widths, we rerun SoFiA-2 with a maximum spectral axis smoothing kernel equal to this apparent width.
@@
Figure 2 presents spectra for each target, binned to lower spectral resolution where sensible333See Table 5 in Appendix for more information.. The GMRT and MeerKAT spectra are generated by summing the flux from pixels included in a 2D mask, which itself is a flattened projection of the 3D mask generated by SoFiA-2. We also present moment 0 (intensity), moment 1 (velocity), and moment 2 (velocity dispersion) maps for these targets in Figure 3, overlaid on imaging from the DESI Legacy Imaging Survey DESI (25), the Dark Energy Camera Plane Survey (DECaPS, 78, DECaPS), and the Pan-STARRs1 survey PanSTARRS (13). The lowest contours are defined as the 3 level measured by SoFiA-2††footnotemark: .
@@
Table 3 presents a summary of the global Hi properties of the sample. These properties are measured from the masked spectra; i.e. only voxels in the 3D SoFiA-2 mask are included, and as a result they are slightly different to those displayed in Fig. 2. We calculate Hi masses following (Meyer2017, 65, equation 48 ), assuming a 10% absolute flux scale uncertainty on all measurements due to calibration alongside the noise contribution. As radio continuum (RC) emission is considered a robust tracer of star formation due to its association with young, massive stars, when RC sources are present, we estimate an SFR following the methods of (Molnar2021, 67, equation 22 ).
@@
We estimate the radius of the Hi emission, HI, out to the 3 contour level by measuring the average surface density33footnotemark: 3 of the intensity maps as a function of elliptical radius from a chosen centre, which we choose to be the optical centre of brightness in the survey imaging. This HI estimate is technically a lower limit; as many of our maps do not reach the nominal 1 pc-2 threshold used in the literature, they are not necessarily half the disk size commonly used to describe HI disks e.g. for the mass-size relation Wang2016 (87). We estimate the ellipticity and position angles of the galaxies using 2MASS 2MASS (83) K-band data where possible; otherwise we fit elliptical isophotes to the legacy photometry presented in Fig. 3. These values are then combined with estimates of the galaxies’ rotation velocities to make broad approximations of the dynamical mass and hence the virial mass for these galaxies. We estimate using
@@
where is the line width at 50% of the peak flux and is the inclination of the galaxy; we discard four galaxies with inclinations lower than 10. To ensure these rotational velocities are in the flat part of each galaxies’ rotation curve, we examine position-velocity diagrams across the major axes and find that all targets extend beyond the break. We do not apply systematic corrections for instrumental broadening and turbulence as we expect such corrections to be small for such high resolution data, and we are not attempting a fully robust measurement. We calculate and following (Yu2020, 92, equations 10 & 13 ).
@@ 3.1 Marginal / Non-Detections
As seen in Fig. 2, the host of FRB 20201123A peaks very marginally above the noise. This host is at a relatively high redshift of 0.0507, and the MeerKAT observation was 0.4 from the phase centre. Due to this, and the beam size, we presume the picked-up signal is simply the unresolved core emission; as such, we do not place much weight on its derived Hi properties, considering its Hi mass to be a lower limit. Only its intensity map is displayed in Fig. 8, and we do not attempt to measure its asymmetry.
Fig. 2 also shows the non-detection of the host of FRB 20200223B. We place a 5 upper limit on its Hi mass using both the GMRT and FAST data, assuming a linewidth of 200 km s-1 and a disk size of 30 kpc. As outlined in Ibik2023 (43), this host is in a transitional regime between star-formation and quiescence, which reflects its lower gas fraction. This is uncommon for FRB hosts; Gordon2023 (34, 81) clearly find a preference towards star-forming galaxies, a conclusion well supported by our findings given that 16 of our 17 targets are emitting in Hi.
@@ 4.1 Asymmetry Metrics
There are a number of common metrics used in the literature to assess asymmetry (Deg2023, 24, see the introduction of); in this study we utilise five: flux, spec, 1D, , and 3D. The first is a common historical measurement used to consider the lopsidedness of a spectrum, and is defined by
@@
The second metric, spec, was introduced by Reynolds2020LVHIS (75) to emphasise more local disturbances through investigation of channel-by-channel variation, defined by
where and are the fluxes of the channel and its mirrored channel pair about a central channel. This centre is generally different than , and is instead chosen as the flux-weighted mean velocity to highlight the differences in profile far away from the centre of mass, e.g. through an extended tail on one side of the spectrum. These metrics were used in Michalowski21 (66, 31) to compare FRB hosts with wider LVHIS and HALOGAS Hi survey populations Reynolds2020LVHIS (75), and thus we include calculations here.
Recently, Deg2023 (24) introduced new metrics to properly account for the bias contributed by noise to the asymmetry of a profile, and to take advantage of the full information that three-dimensional datacubes can offer. As next generation surveys, such as WALLABY and LADUMA, will resolve many Hi galaxies, these higher dimensional metrics are likely to be the most robust way to quantify a galaxy’s disturbance, especially in low SNR regimes. Thus, we also calculate values for 1D, , and 3D, based on equation 20 in Deg2023 (24):
@@
When extracting final values for each of the 1D metrics, we consider the optimal spectral resolution. Each can become temperamental as the SNR of the Hi line decreases and as the linewidth (in channels) decreases, and thus we take measurements at many different rebinning ratios (i.e. until the spectral line is unresolved) to estimate the most stable resolution. This estimation differs between target and is done purely by eye; we inspect the variation of each metric to find the resolution which appears coincident with the most overall stable point in the three metrics, aiming to use the highest resolution possible. We highlight an example case in Fig. 16 in the Appendix.
@@ 4.2 Asymmetry Thresholds
We wish to identify the thresholds above which a profile is considered asymmetric for each of our metrics. Espada2011 (29) modelled the flux distribution of AMIGA isolated galaxies, finding a 2 threshold of 1.26. Reynolds2020LVHIS (75) did the same with the low density fields of LVHIS and HALOGAS, and the higher density field of VIVA, finding 2 thresholds of 1.25, 1.17, and 1.53 respectively for galaxies with stellar masses . Watts2020 (88) focused on the more representative xGASS sample, and chose to define asymmetry by generating noiseless mock spectra with inherent flux equal to 1.1, perturbing them by various noise levels, and finding the threshold (as a function of SNR) within which 80% of the spectra are measured. Considering these results and the fact that a number of our targets are low SNR (10), we choose a flat threshold of 1.3; this is similar to the 3 limits of Espada2011 (29, 75), and corresponds to the low SNR threshold of Watts2020 (88).
Reynolds2020LVHIS (75) also measured spec for the LVHIS, HALOGAS, and VIVA samples, finding 2 thresholds of 0.428, 0.263, and 2.031 respectively. As this metric is the most affected by noise, we choose a slightly arbitrary conservative threshold of 0.5; this also reflects the fact that FRB hosts are not exclusively found in low-density environments.
As the 1D, , and 3D measurements are even newer, there are not many comparisons to be made with wider populations from which to infer at what point a profile is considered asymmetric. Perron-Cormier2025 (69) recently applied these metrics to data from the WALLABY pilot survey Westmeier2022 (90) and found that an 3D threshold of 0.5 corresponds to highly disturbed galaxies, though this value likely represents the extreme end of the disturbance spectrum. Fig. 8 of Deg2023 (24) presents a trial analysis of 500 galaxies from a simulated WALLABY-like survey, showing a median visual disturbance of approximately 0.2, with most values falling within 0.1 of this median. Thus, an 3D value of 0.3 can serve as a rough threshold for identifying galaxies with above-average disturbance. In Section 5, we also present asymmetry measurements of the disturbed FRB host from Lee-waddell23 (52) to use as a reference.